Exactly 12 minutes after the Smiths head north on the turnpike, the Jones set out from the same point to overtake them. The smiths travel at a steady 45 mph, the Jones travel at 54 mph. How long does it take the Jones to overtake the Smiths?
Answer by mananth(16949)

Smiths 45 mph
Jones 54 mph
Difference in time= 00:12 => 0.2 hours (0.2*45)=9
Smiths will have covered 9 miles before Jones starts
catch up distance= 9 miles
catch up speed = 54 -45 mph
catch up speed = 9 mph
Catchup time = catchup distance/catch up speed
catch up time= 9 / 9
catch up time= 1 hour
m.ananth@hotmail.ca